# Regarding the extension of metaplectic geometrical optics to modelling evanescent waves in ray-tracing codes

### v2

@@ Abstract
Metaplectic geometrical optics (MGO) is a recently developed ray-tracing framework to accurately compute the wavefield behavior near a caustic (turning point or focal point), where traditional ray-tracing breaks down. However, MGO has thus far been restricted to having real-valued wavevectors. This is disadvantageous because often upon crossing a caustic from the ‘illuminated’ region to the ‘shadow’ region, two real-valued rays coalesce into one complex-valued ray corresponding to the transition from propagating to evanescent behavior. One can distinguish caustics as having either ‘illuminated shadows’ or ‘proper shadows’ – the former corresponds to when the shadow still contains real-valued rays (albeit in a fewer quantity than in the illuminated region), while the latter corresponds to when the shadow contains no real-valued rays. Here, by means of examples, we show how MGO can be used to model both types of shadows. First, for illuminated shadows we show that MGO can actually be used ‘as is’, provided a corrected quadrature-angle bias is used compared to that proposed in the original references. This is then implemented and demonstrated in a recently developed MGO ray-tracing code. Second, we show that for proper shadows, the MGO formalism can still be used if the symplectic rotation matrix that removes caustics along rays is allowed to be complex-valued. In both cases, strong agreement is seen between the MGO and the exact solution, demonstrating the potential of MGO for improving the predictive capability of ray-tracing codes and laying the foundations for modeling more complicated evanescent phenomena such as tunnelling with MGO.
@@ I Introduction
Here we undergo two case studies to develop understanding for how MGO can be used to model evanescent waves. First, we study a one-dimensional analogue of a cusp caustic. (The cusp caustic commonly occurs near focal points.) This example is chosen because the exact solution contains the superposition of one propagating wave (corresponding to a real-valued ray) and one evanescent wave (corresponding to a complex-valued ray). We show that, although the complex ray is not formally included within MGO, in practice it is possible to obtain its contribution in a numerical implementation because once the complex ray becomes close enough to the real line (meaning the magnitude of the evanescent wave becomes comparable to the propagating one), it becomes numerically degenerate with the real-valued ray. This principle is demonstrated by showing that a recently developed MGO code Hojlund et al. (2024) can be brought into strong agreement with the exact solution, even though no complex rays are formally included. As a prerequisite result to this, we also correct an error in the quadrature-angle bias proposed originally in Ref. Donnelly et al., 2021, and show that the corrected angle bias enables less-supervised simulations.
@@
This paper is organized as follows: in Sec. II the basic equations of MGO are reviewed; in Sec. III it is demonstrated how MGO can be used to model the ‘illuminated shadow’ of a cusp caustic to reasonable accuracy even though no complex rays are formally included; in Sec. IV a new MGO formalism is proposed to model evanescent waves and is validated by reproducing the exponential decay of the Airy function to remarkably high accuracy; finally, in Sec. V the main results are summarized. Appendices present auxiliary results, namely, a simplified MGO formalism (Appendix A) and the corrected quadrature-angle bias (Appendix B).
@@ II Review of MGO
Lastly, $\mathcal{C}_{0}$ denotes the steepest-descent contour that passes through the saddlepoint at $\epsilon=0$. More is said about $\mathcal{C}_{0}$ in Appendix B, which also describes the corrected quadrature-angle bias for numerical implementations of MGO.
On this note, it is worth remarking that the new angle bias described in Appendix B allows complex saddlepoints (i.e., complex rays) to be included in the MGO solution (3b) when they become most important, even though they are formally absent. Complex saddlepoints become most important when they begin to coalesce with real saddlepoints, otherwise they are exponentially damped Wright (1980). Although formally this coalescence only occurs exactly at a caustic, numerically it will occur in a small neighborhood of the caustic. When that happens, the new angle bias is able to include this additional contribution by effectively integrating along the union of the two individual steepest-descent contours. This is seen in Fig. 5 for the cusp caustic: the steepest-descent valleys tracked by the new angle bias actually correspond to the valleys of the nearby complex saddlepoint, not the real saddlepoint at $\epsilon=0$. When the two saddlepoints become sufficiently close [determined by details of the root-finding procedure (34)], the new algorithm treats them as effectively coalesced. The result is a smoother field than one would obtain if the complex saddlepoints were omitted entirely.
@@ III.2 Numerical results
Now let us evaluate the MGO solution (14) numerically using the Gauss–Freud quadrature rule with the corrected quadrature-angle bias (Appendix B). The computation is performed from $x=-10$ to $x=10$ and crosses over the caustic at $x=0$. No attempt is made to prescribe the caustic location at $x=0$ or impose that the solution be symmetric about $x=0$.
@@
Figure 2: Comparison of the MGO solution for the 1D cusp caustic given by Eq. (14), as computed using the corrected angle bias within the Gauss–Freud quadrature routine [Appendix B], with the standard GO solution given by Eq. (10) and the exact solution given by Eq. (6). A tenth-order quadrature rule is used.
Figure 3: Comparison of the exact solution for the 1D cusp caustic given by Eq. (6) with the solution provided by the MGO ray-tracing code developed in Ref. Hojlund et al., 2024 after the new angle bias has been implemented in the Gauss–Freud quadrature routine. Shown in red is the raw, unfiltered output from the MGO code, while shown in blue is the result of applying additional high-$k$ filtering to the code output. Specifically, all fluctuations with wavevectors $k>10k_{\text{max}}$ are removed, where $k_{\text{max}}=10^{1/3}$ is the maximum wavevector expected in the solution.
The corrected quadrature-angle bias has also been implemented in the MGO code developed in Ref. Hojlund et al., 2024, which is a complete numerical implementation of MGO. [This means that the exact expressions that comprise Eq. (14) are never entered anywhere as input, but are instead computed from first principles.] The result of applying this code (with corrected angle bias) to the cusp differential equation (5) is shown in Fig. 3. The raw output of the code (shown in transparent red) is relatively well-behaved; the residual errors are due to spurious poles (Froissart doublets) that get introduced due to the numerical analytical-continuation procedure Hojlund et al. (2024) and might be removable by applying a more advanced version of the analytical-continuation algorithm Nakatsukasa et al. (2018).
A simpler remedy is to apply a low-pass filter based on the maximum wavevector expected in the solution. Indeed, since the local wavevector of the solution is already known from the ray trajectory $k(\tau)$, any oscillations with wavelengths much larger than $\text{max}[k(\tau)]$ are spurious and should be filtered out. The results of applying such a low-pass filter to the code output is also shown in Fig. 3, with visible improvements. It is interesting to note that with the original quadrature-angle bias, the fluctuations in the raw code output were so great that even applying a low-pass filter could not get the solution to be well-behaved near the caustic.
@@ IV Modeling ‘proper shadows’ with complex MGO: Fold caustic case study
To generalize Eq. (3b) for modeling complex-ray contributions with MGO when there are no real-valued rays, we propose to use the following formula based on the general derivation of MGO presented in Ref. Lopez, 2022: [^5] Specifically, one defines the symplectic rotation matrix to now be complex-valued and unitary. This imposes that the matrix sub-blocks $A$, $B$, $C$, and $D$ satisfy $A=D^{*}$, $C=-B^{*}$, and $|D|^{2}+|C|^{2}=1$. Finally, as outlined in Ref. Lopez, 2022, one then assigns $D=\dot{x}/|\dot{{\boldsymbol{\rm z}}}|$ and $C=-\dot{k}/|\dot{{\boldsymbol{\rm z}}}|$, where ${\boldsymbol{\rm z}}=(x,k)$.
@@ V Conclusion
The second case study is the fold caustic, whose shadow consists of only a single complex ray corresponding to a purely decaying wavefield. We propose a simple complexification of MGO in which the formalism remains the same but the constituent expressions are allowed to be complex. The resulting complex-MGO solution to Airy’s equation shows remarkable agreement with the exact result, with no visible difference between the two even when viewed on a logarithmic scale.

### v3

@@ Abstract
Metaplectic geometrical optics (MGO) is a recently developed ray-tracing framework to accurately compute the wavefield behavior near a caustic (turning point or focal point), where traditional ray-tracing breaks down. However, MGO has thus far been restricted to having real-valued wavevectors. This is disadvantageous because often upon crossing a caustic from the ‘illuminated’ region to the ‘shadow’ region, two real-valued rays coalesce into one complex-valued ray corresponding to the transition from propagating to evanescent behavior. One can distinguish caustics as having either ‘illuminated shadows’ or ‘proper shadows’ – the former corresponds to when the shadow still contains real-valued rays (albeit in a fewer quantity than in the illuminated region), while the latter corresponds to when the shadow contains no real-valued rays. Here, by means of examples, we show how MGO can be used to model both types of shadows. First, for illuminated shadows we show that MGO can actually be used ‘as is’, provided a corrected integration scheme is used compared to that proposed in the original references. This is then implemented and demonstrated in a recently developed MGO ray-tracing code. Second, we show that for proper shadows, the MGO formalism can still be used if the symplectic rotation matrix that removes caustics along rays is allowed to be complex-valued. In both cases, strong agreement is seen between the MGO and the exact solution, demonstrating the potential of MGO for improving the predictive capability of ray-tracing codes and laying the foundations for modeling more complicated evanescent phenomena such as tunnelling with MGO.
@@ I Introduction
Here we undergo two case studies to develop understanding for how MGO can be used to model evanescent waves. First, we study a one-dimensional analogue of a cusp caustic. (The cusp caustic commonly occurs near focal points.) This example is chosen because the exact solution contains the superposition of one propagating wave (corresponding to a real-valued ray) and one evanescent wave (corresponding to a complex-valued ray). We show that, although the complex ray is not formally included within MGO, in practice it is possible to obtain its contribution in a numerical implementation because once the complex ray becomes close enough to the real line (meaning the magnitude of the evanescent wave becomes comparable to the propagating one), it becomes numerically degenerate with the real-valued ray. This principle is demonstrated by showing that a recently developed MGO code Hojlund et al. (2024) can be brought into strong agreement with the exact solution, even though no complex rays are formally included. As a prerequisite result to this, we also correct an error in the quadrature rule proposed originally in Ref. Donnelly et al., 2021, and show that the corrected method enables less-supervised simulations.
@@
This paper is organized as follows: in Sec. II the basic equations of MGO are reviewed; in Sec. III it is demonstrated how MGO can be used to model the ‘illuminated shadow’ of a cusp caustic to reasonable accuracy even though no complex rays are formally included; in Sec. IV a new MGO formalism is proposed to model evanescent waves and is validated by reproducing the exponential decay of the Airy function to remarkably high accuracy; finally, in Sec. V the main results are summarized. Appendices present auxiliary results, namely, a simplified MGO formalism (Appendix A) and the corrected quadrature rule (Appendix B).
@@ II Review of MGO
Lastly, $\mathcal{C}_{0}$ denotes the steepest-descent contour that passes through the saddlepoint at $\epsilon=0$. More is said about $\mathcal{C}_{0}$ in Appendix B, which also describes the corrected quadrature rule for numerical implementations of MGO. In Eqs. (II) and (4), all of $x$, $k$, and $\tau$ are real-valued, but this constraint will be re-evaluated in the following sections.
On this note, it is worth remarking that the new quadrature rule described in Appendix B allows complex saddlepoints (i.e., complex rays) to be included in the MGO solution (3b) when they become most important, even though they are formally absent. Complex saddlepoints become most important when they begin to coalesce with real saddlepoints, otherwise they are exponentially damped Wright (1980). Although formally this coalescence only occurs exactly at a caustic, numerically it will occur in a small neighborhood of the caustic. When that happens, the new method is able to include this additional contribution by effectively integrating along the union of the two individual steepest-descent contours. This is seen in Fig. 5 for the cusp caustic: the steepest-descent valleys tracked by the new method actually correspond to the valleys of the nearby complex saddlepoint, not the real saddlepoint at $\epsilon=0$. When the two saddlepoints become sufficiently close [determined by details of the root-finding procedure (34)], the new algorithm treats them as effectively coalesced. The result is a smoother field than one would obtain if the complex saddlepoints were omitted entirely.
@@ III.2 Numerical results
Now let us evaluate the MGO solution (14) numerically using a tenth-order Gauss–Freud quadrature rule (modified according to Appendix B). The computation is performed from $x=-10$ to $x=10$ and crosses over the caustic at $x=0$. No attempt is made to prescribe the caustic location at $x=0$ or impose that the solution be symmetric about $x=0$.
@@
Figure 2: Comparison of the MGO solution for the 1D cusp caustic given by Eq. (14) with the standard GO solution given by Eq. (10) and the exact solution given by Eq. (6).
Figure 3: Comparison of the exact solution for the 1D cusp caustic given by Eq. (6) with the solution provided by the MGO ray-tracing code developed in Ref. Hojlund et al., 2024 after the new quadrature rule has been implemented. Shown in red is the raw, unfiltered output from the MGO code, while shown in blue is the result of applying additional high-$k$ filtering to the code output. Specifically, all fluctuations with wavevectors $k>10k_{\text{max}}$ are removed, where $k_{\text{max}}=10^{1/3}$ is the maximum wavevector expected in the solution.
The corrected quadrature rule has also been implemented in the MGO code developed in Ref. Hojlund et al., 2024, which is a complete numerical implementation of MGO. [This means that the exact expressions that comprise Eq. (14) are never entered anywhere as input, but are instead computed from first principles.] The result of applying this code to the cusp differential equation (5) is shown in Fig. 3. The raw output of the code (shown in transparent red) is relatively well-behaved; the residual errors are due to spurious poles (Froissart doublets) that get introduced due to the numerical analytical-continuation procedure Hojlund et al. (2024) and might be removable by applying a more advanced version of the analytical-continuation algorithm Nakatsukasa et al. (2018).
A simpler remedy is to apply a low-pass filter based on the maximum wavevector expected in the solution. Indeed, since the local wavevector of the solution is already known from the ray trajectory $k(\tau)$, any oscillations with wavevectors much larger than $\text{max}[k(\tau)]$ are spurious and should be filtered out. The results of applying such a low-pass filter to the code output is also shown in Fig. 3, with visible improvements. It is interesting to note that with the original quadrature rule, the fluctuations in the raw code output were so great that even applying a low-pass filter could not get the solution to be well-behaved near the caustic.
@@ IV Modeling ‘proper shadows’ with complex MGO: Fold caustic case study
To generalize Eq. (3b) for modeling complex-ray contributions with MGO when there are no real-valued rays, we propose to use the following formula based on the general derivation of MGO presented in Ref. Lopez, 2022: [^5] Specifically, one repeats the derivation of MGO presented in Ref. Lopez, 2022 but now with the underlying $2\times 2$ symplectic matrix $S$ being complex-valued and unitary; the matrix elements are $S_{22}=S_{11}^{*}=\dot{x}/|\dot{{\boldsymbol{\rm z}}}|$ and $S_{21}=-S_{12}^{*}=-\dot{k}/|\dot{{\boldsymbol{\rm z}}}|$, with ${\boldsymbol{\rm z}}=(x,k)$.
@@ V Conclusion
The second case study is the fold caustic, whose shadow consists of only a single complex ray corresponding to a purely decaying wavefield. We propose a simple complexification of MGO in which the formalism remains the same but the constituent expressions are allowed to be complex. The resulting complex-MGO solution (26) to Airy’s equation shows remarkable agreement with the exact result, with no visible difference between the two even when viewed on a logarithmic scale. This formula can be used in MGO codes to include evanescent fields via local asymptotic matching.
@@ Acknowledgements
Disclosures
The authors declare no conflicts of interest.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.